# SVG360: Multi-View SVG Generation with Geometric and Color Consistency from a Single SVG

### v1

@@ 3.2 Spatially Aligned Segmentation Propagation
After appearance harmonization, the multi-view rasters exhibit more uniform regions and sharper boundaries. To obtain cleaner and semantically stable vector paths in the subsequent vectorization stage, we need to maintain consistency of part-level identity across views. Relying on single view segmentation alone cannot enforce cross-view alignment of labels and boundaries, which weakens the robustness of path fitting and cross-view correspondence. Multi-view SVG generation inherently requires segmentation to be continuous in space. Accordingly, we spatialize the continuous segmentation paradigm of SAM2 by replacing temporal adjacency with spatial neighborhood on the viewing sphere and propagating masks along that sphere. Each target view is segmented under guidance from its nearest spatial neighbors, thereby maintaining part identity and boundary consistency at a global level.
Spatially Sequential Scheduling. We uniformly sample camera viewpoints on a unit sphere to obtain a globally distributed set of directions covering the entire view space. The sampling density and angular step are adjustable, allowing flexible control over the number of viewpoints and their spatial distribution according to task requirements. This configuration achieves globally balanced coverage even with a limited number of samples, forming a stable spatial adjacency structure that serves as the geometric foundation for subsequent cross-view propagation and visibility-aware reprojection. Although the sampled views have no intrinsic temporal order, a pseudo-sequential traversal helps maintain continuity during segmentation propagation. Starting from the front view aligned with the input SVG (denoted as $\theta_{0}$), each subsequent view is chosen as the unvisited one with the smallest angular distance from the current view, until all views are processed:
@@
Reference View Selection. After determining the traversal order, we construct a reference set for each target view to reuse information from geometrically adjacent and previously processed views. For each target view $\theta$, we select its $k$ nearest processed neighbors within an angular threshold $\tau$, ensuring that the guidance originates from spatially reliable local contexts. We empirically set $k=6$ and $\tau=75^{\circ}$ in all experiments. This localized reference selection, combined with the unified angular distance metric, restricts soft prompting and mask propagation to geometrically adjacent views, thereby preserving consistent part identities and boundary alignment across the entire viewing sphere.
Key Frame Initialization and Filtering.
Segmentation starts from the key view $\theta_{0}$, which corresponds to the input SVG. After obtaining initial masks from the automatic segmentation model, we apply a lightweight post-processing to reduce redundancy and improve stability for subsequent propagation. Small regions smaller than either 200 pixels or 0.05% of the image area (whichever is larger) are removed, a single morphological closing with an elliptical kernel sized to about 0.5% of the shorter image dimension is used to smooth boundaries, and overlapping masks are suppressed using non-maximum suppression with an IoU threshold of 0.5. The cleaned masks serve as prompt inputs to the spatially SAM2, providing stable guidance for the following multi-view segmentation process.
Residual Discovery.
Since the initial segmentation relies solely on the key view $\theta_{0}$, previously unseen parts that appear in other viewpoints may remain undetected. After Spatial-SAM2 completes segmentation across all sampled views, we introduce a residual discovery loop to automatically detect and recover these missing regions. For each processed view $\theta_{t}$, we compute the uncovered foreground ratio $r$ based on the union of all current masks. If $r$ exceeds 5% of the estimated foreground area, the same automatic segmentation procedure used for the key view is reactivated, but restricted to the uncovered regions. Newly discovered masks are then merged into the existing mask set and propagated to subsequent views through Spatial SAM2 to maintain cross-view consistency. The overall process is summarized in Algorithm 1.
@@
        Propagate updates to neighboring views via Spatial-SAM2;
@@ 3.3 Vector Consolidation
Figure 2: Qualitative comparison. The figure summarizes representative issues observed in Adobe Turntable: (a) geometric inconsistencies across adjacent views; (b) cluttered structures arising from overlapping thin components; (c) merging of parts with similar colors; (d) missing regions in certain viewpoints; (e) color drift and gradual loss of small details across views. Our method produces multi-view SVGs with stable geometry, clear part separation, and consistent color appearance.
After segmentation, we crop each part into a transparent raster and feed it to VTracer [37], which converts pixel regions into path sets by color layering and spline fitting. Processing per part rather than per image reduces cross-part interference and yields a more compact path topology. This simplifies later consolidation and improves cross-view correspondence.
@@
To reduce small cross-view color drift, we extract a reference palette from the input SVG and map each source color to the nearest palette entry using the CIEDE2000 color-difference metric [31]. We add a near-black bias to stabilize dark tones: colors in low lightness and low chroma preferentially map to black, otherwise to the closest non-black reference color. This preserves semantic colors while suppressing banding due to shadows and reflections. In practice we preserve alpha values, resolve inherited attributes into explicit colors, and remove empty attributes so that rendering is consistent across editors. Together, color alignment unifies appearance and consolidation compresses structure, improving cross-view consistency and the editability of the final SVGs.
@@ 4.1 Implementation Details
For the 3D generation stage, we employ the TRELLIS-image-large model to generate multi-view rasterizations from the input SVG. To harmonize the rendered appearance, lightweight LoRA adapter is trained to adapt FLUX.1-dev to SVG-style appearance, following standard PEFT settings (rank=32). The text encoder is frozen, and the adapter is trained for 1,000 steps. All experiments are conducted on a single NVIDIA A100-80GB PCIe GPU.
@@ 4.2 Evaluation Metrics
We use two metrics to capture cross-view stability. $\mathrm{RMSE}_{\text{path}}$ measures how much each view’s path count deviates from the input SVG. $\overline{\Delta N}_{\text{color,nbr}}$ measures the change in color count between adjacent views, lower indicates more stable color behavior.
@@ 4.3 Comparison
Baseline Adobe Illustrator Turntable (Beta) is a commercial, publicly available end-to-end feature that generates multi-view SVGs from a single 2D vector artwork. Its internal implementation, view grid, and technical specifications are not disclosed, and no programmatic interface is provided. We therefore use Turntable as the industrial baseline for our full pipeline comparison.
According to the official documentation, Turntable supports approximately $\pm 120^{\circ}$ horizontal rotation in yaw (sampled every $15^{\circ}$) and three pitch levels at $\{-45^{\circ},0^{\circ},+45^{\circ}\}$. For fair comparison, we adopt the same camera poses defined by the Turntable configuration in our experiments.
Because Turntable only allows manual per-view SVG export, we evaluate this baseline on 10 representative SVG examples from which complete view sequences can be consistently obtained.
@@
Quantitative Comparison Table 1 reports the numerical statistics for Adobe Turntable and our method. Our approach produces fewer paths and colors per view, indicating cleaner vector topology and a more compact palette. For cross-view stability, our method achieves lower $\mathrm{RMSE}_{\text{path}}$ and a substantially lower $\overline{\Delta N}_{\text{color,nbr}}$, showing that the generated views remain closer to the original SVG structure and maintain more consistent color behavior across viewpoints. Together, these results confirm that our multi-view SVGs are structurally simpler and exhibit smoother appearance transitions.
Table 1: Comparison of multi-view SVG statistics between Adobe Turntable and our method.
Metric Adobe Turntable Ours
$\overline{N}_{\text{path}}$ 29.59 26.15
$\overline{N}_{\text{color}}$ 4.80 3.37
@@
$\mathrm{RMSE}_{\text{path}}$  ( $\downarrow$ ) 16.37 12.03
$\overline{\Delta N}_{\text{color,nbr}}$  ( $\downarrow$ ) 2.12 0.35
Qualitative Comparison. Figure 2 summarizes the typical failure patterns we observe when applying Adobe Turntable to single-SVG multi-view generation. Across multiple objects and views, Adobe Turntable frequently exhibits geometric inconsistencies, structural confusion, unstable camera poses, view-dependent clipping, and noticeable color drift. In contrast, our method maintains coherent geometry, clear part separation, and consistent appearance across views.
In terms of geometry, object structures do not remain consistent across views. In example (a), components that should be rigid, such as the roof rack and the headlights—change noticeably between yaw $0^{\circ}$ and $90^{\circ}$, with the rack switching from two crossbars to a flat plate and the paired headlights collapsing into a single central light. Examples (b) and (c) further show that when the input contains multiple thin or closely spaced parts, the generated views often introduce overlaps, clutter, or unintended merging, making individual components difficult to distinguish.
Viewpoint-related inconsistencies are also apparent. In examples (a) and (e), the result at yaw $90^{\circ}$ deviates from a clean side view, exhibiting a slight tilt and geometric deformation. When the input SVG spans a large spatial extent, Turntable does not normalize its scale, causing portions of the object to be clipped outside the rendering canvas, as observed in both (a) and (e).
Color stability across views can also be unreliable. In example (e), the sesame seeds on the burger change brightness across viewpoints, and those on the far side gradually diminish as the camera rotates. In example (c), regions with similar color are merged into a single path, reducing structural separation.
In contrast, our method relies on explicit 3D representation, controlled camera poses, and consistent color and path processing. As a result, the generated multi-view SVGs maintain stable geometry, clear part separation, and coherent appearance across all viewpoints.
@@ 4.4 Ablation Study
Figure 3: Segmentation comparison of Spatial-SAM2 and the original SAM2 tracking mode at view 47. The bottom row shows the each six reference views used for memory support, with view 0 serving as the initial key frame. Results shown here correspond to the first iteration before any subsequent refinement.
@@
Figure 4: Ablation study of four segmentation strategies: Spatial SAM2 (ours), SAM2 Tracking Mode, SAM2 Auto Mode, and a segmentation free VTracer baseline.
As shown in Figure 4, we further compare the impact of four segmentation strategies on the final SVG outputs. Spatial-SAM2 produces the most stable and coherent results. The tracking mode of SAM2 still exhibits occasional misses in certain viewpoints. The Auto mode (i.e., SAM2’s Segment Everything) tends to over-fragment continuous structures due to the lack of cross-view constraints, resulting in inconsistent topology across views. When segmentation is entirely omitted, many parts fail to survive the vectorization stage and cannot be reliably reconstructed, leaving only a few coarse contours. These findings confirm that enforcing spatially consistent segmentation is critical for producing high-quality multi-view SVGs.

### v2

@@ 3.2 Spatially Aligned Segmentation Propagation
After appearance harmonization, the multi-view rasters exhibit more uniform regions and sharper boundaries. To obtain cleaner and semantically stable vector paths in the subsequent vectorization stage, we need to maintain consistency of part-level identity across views. Relying on single-view segmentation alone cannot enforce cross-view alignment of labels and boundaries, which weakens the robustness of path fitting and cross-view correspondence. Multi-view SVG generation inherently requires segmentation to be continuous in space. Accordingly, we spatialize the continuous segmentation paradigm of SAM2 by replacing temporal adjacency with spatial neighborhood on the viewing sphere and propagating masks along that sphere. Each target view is segmented under guidance from its nearest spatial neighbors, thereby maintaining part identity and boundary consistency at a global level.
Spatially Sequential Scheduling
We uniformly sample camera viewpoints on a unit sphere to obtain a globally distributed set of directions covering the entire view space. The sampling density and angular step are adjustable, allowing flexible control over the number of viewpoints and their spatial distribution according to task requirements. This configuration achieves globally balanced coverage even with a limited number of samples, forming a stable spatial adjacency structure that serves as the geometric foundation for subsequent cross-view propagation and visibility-aware reprojection.
Although the sampled views have no intrinsic temporal order, a pseudo-sequential traversal helps maintain continuity during segmentation propagation. Starting from the front view aligned with the input SVG (denoted as $\theta_{0}$), each subsequent view is chosen as the unvisited one with the smallest angular distance from the current view, until all views are processed:
@@
Spatial Memory Selector
After determining the traversal order, we construct a reference set for each target view to reuse information from geometrically adjacent and previously processed views. For each target view $\theta$, we select its $k$ nearest processed neighbors within an angular threshold $\tau$, ensuring that the guidance originates from spatially reliable local contexts. We empirically set $k=6$ and $\tau=75^{\circ}$ in all experiments. This Spatial Memory Selector mechanism, combined with the unified angular distance metric, restricts soft prompting and mask propagation to geometrically adjacent views, thereby preserving consistent part identities and boundary alignment across the entire viewing sphere.
Key Frame Initialization and Filtering. Segmentation starts from the key view $\theta_{0}$, which corresponds to the input SVG. After obtaining initial masks from the automatic segmentation model, we apply lightweight postprocessing to reduce redundancy and improve stability for subsequent propagation. We remove regions whose area is smaller than the larger of 100 pixels and 0.05% of the image area. We then apply a single morphological closing with an elliptical kernel whose size is about 0.5% of the shorter image dimension to smooth boundaries. Finally, we suppress overlapping masks using non-maximum suppression with an IoU threshold of 0.5. The cleaned masks serve as prompts for Spatial SAM2, providing stable guidance for the subsequent multi-view segmentation process.
Residual Discovery
Since the initial segmentation relies solely on the key view $\theta_{0}$, previously unseen parts that appear in other viewpoints may remain undetected. After Spatial SAM2 completes segmentation across all sampled views, we introduce a residual discovery loop to automatically detect and recover these missing regions. For each processed view $\theta_{t}$, we compute the uncovered foreground ratio $r$ based on the union of all current masks. If $r$ exceeds 5% of the estimated foreground area, the same automatic segmentation procedure used for the key view is reactivated, but restricted to the uncovered regions. Newly discovered masks are then merged into the existing mask set and propagated to subsequent views through Spatial SAM2 to maintain cross-view consistency. The overall process is summarized in Algorithm 1.
@@
        Propagate updates to neighboring views via Spatial SAM2;
@@ 3.3 Vector Consolidation
After segmentation, we crop each part into a transparent raster and feed it to VTracer [vtracer2023], which converts pixel regions into path sets by color layering and spline fitting. Processing per part rather than per image reduces cross-part interference and yields a more compact path topology. This simplifies later consolidation and improves cross-view correspondence.
@@
To reduce small cross-view color drift, we extract a reference palette from the input SVG and map each source color to the nearest palette entry using the CIEDE2000 color-difference metric $\Delta E_{2000}$ [sharma2005ciede2000]. We add a near-black bias to stabilize dark tones: colors in low lightness and low chroma preferentially map to black, otherwise to the closest non-black reference color. This preserves semantic colors while suppressing banding due to shadows and reflections. In practice we preserve alpha values, resolve inherited attributes into explicit colors, and remove empty attributes so that rendering is consistent across editors. Together, color alignment unifies appearance and consolidation compresses structure, improving cross-view consistency and the editability of the final SVGs.
@@ 4.1 Implementation Details
For the 3D generation stage, we employ the TRELLIS-image-large model to generate multi-view rasterizations from the input SVG. To harmonize the rendered appearance, we use FLUX.1-dev [blackforestlabs2024flux] with a public vector-style LoRA adapter (SimpleVectorFlux [renderartist2024simplevectorflux, hu2021lora]), without any additional training. We follow the adapter’s recommended trigger tokens (v3ct0r, vector) and keep a fixed prompt template across all views for consistent style. All experiments are conducted on a single NVIDIA A100-80GB PCIe GPU.
@@ 4.2 Evaluation Metrics
To comprehensively evaluate our framework, we group our metrics into three aspects. SVG-Centric Statistics: We evaluate the compactness of generated SVGs using the average number of paths ($\overline{N}_{\text{path}}$) and colors ($\overline{N}_{\text{color}}$). To measure structural and appearance stability relative to the input, we use $\mathrm{RMSE}_{\text{path}}$ for path count deviation and $\overline{\Delta N}_{\text{color,nbr}}$ alongside the CIEDE2000 metric $\Delta E_{2000}$ [sharma2005ciede2000] to capture color count changes and perceptual color drift across views. Multi-View Continuity: For appearance continuity on the yaw-pitch viewing grid, we compute DINOv2 [oquab2023dinov2] feature consistency and DreamSim [fu2023dreamsim] perceptual similarity on adjacent view pairs. Specifically, we form yaw-adjacent pairs by grouping views with the same pitch and consecutive yaw steps, and pitch-adjacent pairs by grouping views with the same yaw and consecutive pitch levels. We report Yaw and Pitch by averaging scores over the corresponding adjacent pairs, and Overall by averaging over all adjacent pairs on the grid. Single-View Fidelity: We use standard image quality metrics, SSIM [wang2004image] and LPIPS [zhang2018unreasonable], to evaluate the fundamental raster-to-SVG reconstruction quality. Detailed mathematical formulations and computation specifics for all aforementioned metrics are provided in the supplementary material.
@@ 4.3 Comparison
Baselines.
For multi-view generation, Adobe Illustrator’s Turntable (Beta) [adobeTurntableHelp] serves as our primary industrial baseline. It is a commercial, publicly available end-to-end feature that generates multi-view SVGs from a single 2D vector artwork. Its internal implementation, view grid, and technical specifications are not disclosed, and no programmatic interface is provided. According to the official documentation, Turntable supports approximately $\pm 120^{\circ}$ horizontal rotation in yaw (sampled every $15^{\circ}$) and three pitch levels at $\{-45^{\circ},0^{\circ},+45^{\circ}\}$. For a fair comparison, we adopt the exact same camera poses defined by the Turntable configuration in our experiments. Since Adobe Turntable lacks an API and necessitates manual per-view export, large-scale evaluation is not practical to scale. To ensure rigorous comparison, we curated a representative benchmark of 30 SVGs across categories like vehicles, buildings, and animals. By fully exporting the 51-view yaw-pitch grid for each, we obtained a dense set of 1530 multi-view SVG outputs for strict quantitative and qualitative evaluation. For single view vectorization, we additionally benchmark our vectorization module on a small set of 200 simple single object SVGs and compare against LIVE [yue2024live], StarVector [wang2024starvector], and AutoTrace [weber2001autotrace].
Figure 2: Qualitative comparison with Adobe Turntable. Target camera poses (yaw and pitch) are annotated at the top; ! marks views that deviate from the target pose. (a) Car: unrealistic structural changes (e.g., incomplete or misplaced headlights) and inconsistent foreshortening/scale (e.g., wheels vs. chassis). (b) Torii gate: slender hanging components suffer from path crowding and tangling in certain views. (c) Scooter: distorted handlebar proportions in oblique views and structural inconsistency around the steering tube (head-tube region). (d) Turtle: missing small parts (e.g., nose) and locally incomplete shell regions in challenging views.
Quantitative Comparison.
Multi-View SVG Statistics. Table 1 compares the intrinsic properties of generated SVGs. Our approach produces highly compact vectors (avg. 20.77 paths, 3.65 colors) while demonstrating exceptional stability across viewpoints. Compared to Adobe Turntable, we reduce path count deviation ($\mathrm{RMSE}_{\text{path}}$) from 14.22 to 10.95. Color stability improvements are particularly striking: we reduce adjacent-view color count variation ($\overline{\Delta N}_{\text{color,nbr}}$) from 2.01 to 0.29, and perceptual color drift ($\Delta E_{2000}$) from 3.22 to 1.19, substantially reducing the structural flickering and color inconsistencies common in the baseline.
@@
Table 1: SVG-centric multi-view comparison against Adobe Turntable. We report statistics of the generated multi-view SVGs, including compactness (fewer paths/colors indicate a more compact and editable SVG) and cross-view stability derived from SVG structure and palette behavior. Lower $\mathrm{RMSE}_{\text{path}}$, $\overline{\Delta N}_{\text{color,nbr}}$, and $\Delta E_{2000}$ indicate more stable structure and color across views.
Method Compactness SVG structure Color consistency
@@
$\overline{N}_{\text{path}}$ $\overline{N}_{\text{color}}$ $\mathrm{RMSE}_{\text{path}}\downarrow$ $\overline{\Delta N}_{\text{color,nbr}}\downarrow$ $\Delta E_{2000}\downarrow$
Adobe Turntable 22.34 4.41 14.22 2.01 3.22
Ours 20.77 3.65 10.95 0.29 1.19
Multi-View Continuity. Table 2 evaluates spatial continuity across the yaw-pitch viewing grid. Overall, our approach yields more coherent view-to-view transitions than Adobe Turntable. While Turntable attains competitive perceptual scores on simple yaw rotations, our method is more robust on the full grid, achieving higher overall DINO consistency (0.8864 vs. 0.8611). Overall DreamSim is comparable (0.8392 vs. 0.8388), but the advantage becomes clear under challenging pitch changes, where our method achieves higher Pitch DreamSim (0.7748 vs. 0.7563), indicating smoother appearance transitions across viewpoints.
Table 2: Multi view continuity comparison against Adobe Turntable on a yaw and pitch camera grid. We measure local transition smoothness between adjacent views using DINOv2 feature consistency and DreamSim perceptual similarity, reporting yaw, pitch, and overall scores.
Method DINO Consistency DreamSim Similarity
[…]
Adobe Turntable 0.9005 0.8053 0.8611 0.8973 0.7563 0.8388
Ours 0.9074 0.8567 0.8864 0.8848 0.7748 0.8392
Single-View Raster-to-SVG Fidelity. Table 3 validates our vectorization module against state-of-the-art models. Our method yields the highest visual fidelity (SSIM: 0.926, LPIPS: 0.020, $\Delta E_{2000}$: 0.408) while striking an optimal balance in structural compactness. Specifically, while StarVector produces slightly fewer paths (15.2) at the cost of poor fidelity ($\Delta E_{2000}$: 17.706), and AutoTrace yields reasonable similarity (SSIM: 0.812) but heavily fragmented topology (448.2 paths), our method avoids these trade-offs, delivering near-perfect reconstruction with only 19.75 paths.
Table 3: Single-view (initial input view) raster-to-SVG comparison. We render predicted SVGs back to rasters and evaluate against the input raster.
[…]
LIVE [yue2024live] 0.803 0.065 7.364 18.2
StarVector [wang2024starvector] 0.697 0.153 17.706 15.2
AutoTrace [weber2001autotrace] 0.812 0.063 2.792 448.2
Ours 0.926 0.020 0.408 19.75
Qualitative Comparison.
Figure 2 provides a visual comparison with Adobe Turntable. Overall, the baseline often struggles to maintain cross-view 3D consistency for single-SVG multi-view generation. It is prone to geometric/foreshortening distortions, cluttered paths on thin or overlapping structures, and occasional pose drift. Under challenging viewpoints, it may also miss small semantic parts. In contrast, by leveraging an explicit 3D representation with strict pose control, our approach maintains more stable geometry, cleaner part separation, and coherent multi-view appearances.
@@ 4.4 Ablation Study
Figure 3: Segmentation comparison between the original SAM2 tracking mode (left) and Spatial SAM2 (ours, right) at view 47. For each method, we show the six reference views used for memory support. Tracking SAM2 uses views 41-46, while Spatial SAM2 uses views 24-26, and 44-46. Results correspond to the first iteration before any subsequent refinement.
@@
As shown in Figure 4, we further compare the impact of four segmentation strategies on the final SVG outputs. Spatial SAM2 produces the most stable and coherent results. The tracking mode of SAM2 still exhibits occasional misses in certain viewpoints. The Auto mode (i.e., SAM2’s Segment Everything) tends to over-fragment continuous structures due to the lack of cross-view constraints, resulting in inconsistent topology across views. When segmentation is entirely omitted, many parts fail to survive the vectorization stage and cannot be reliably reconstructed, leaving only a few coarse contours. These findings confirm that enforcing spatially consistent segmentation is critical for producing high-quality multi-view SVGs.
Figure 4: Ablation study of four segmentation strategies: Spatial SAM2 (ours), SAM2 Tracking Mode, SAM2 Auto Mode, and a segmentation-free VTracer baseline.